This is Just one question not multipple, please help. I always rate. States to remember to write out your calculations and explanations for ALL problems.

Remember to write out your calculations and explanations for ALL problems. As usual, if you nccd help with these problems, you may ask mc for help cither privately through Blackboard email/chat or by calling me, or you may also post your questions directly under the discussion for this project. Alternatively, you may also work on these with a tutor at the Student Academic Support Center on campus. Problem Set 6 Problems on Gene Expression/Repa r and Protein Synthesis/Degradation 3 problems) 16. The DNA in a human non-gametic cell contains 6 billion base pairs. It is estimated that about 10,000 DNA changes occur in each cell in one day. These are quickly repaired so that only a few (1 to 5) mutations accumulate in one cell in a year a. What percentage of the base pairs are altered each day? b. What percentage of the DNA changes that occur in one cell in one year escape the proofreading and repair process calculate this in both cases: if 1 mutation accumulates in one cell per year, and if 5 mutations accumulate in one cell per year?
Answer #1

note that each DNA change refers to one base pair

a)percentage of base pairs affected each day = 10000/1000000000 x 100 = 0.001 % DNA changes per day

b) total number of DNA changes that happen in one year= 10000 x 365 = 36.5 x 104

so if one change escapes from DNA repair mechanisms, its % among 36/5 x 104 will be

1/36.5x104 X 100% = 0.000273926%

same case but if 5 changes escape DNA repair mechanisms ,the % of those 5 among 36.5 x 104 would be

5/36.5x104   X 100% = 0.00136986301%